Two spherical objects have masses of 1.5 x 10^5 kg and 8.5 x 10^2 kg. Their centers are separated by a distance of 2500 m. Find the gravitational attraction between them
1
Expert's answer
2020-12-08T10:51:49-0500

We can find the gravitational attraction between two spherical objects from the Newton's law of universal gravitation:


"F=\dfrac{Gm_1m_2}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot 1.5\cdot10^5\ kg\cdot 8.5\cdot10^2\ kg}{(2500\ m)^2}=1.36\cdot10^{-9}\ N."

Answer:

"F=1.36\cdot10^{-9}\ N."